OpenStudy (anonymous):

FAN AND MEDAL: Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x - 1)? A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 greater than or equal to x x less than or greater to 3

OpenStudy (anonymous):

a number line with a closed circle on 3 and shading to the right

OpenStudy (anonymous):

Thank You! @Wolverine32 Can you help me with a few more?

OpenStudy (anonymous):

@ranga @Hero @PinkSapphire @PINKYPIE19904 @Jagking

OpenStudy (anonymous):

Shiloh has to earn at least $200 to meet her fundraising goal. She has only 100 cakes that she plans to sell at $5 each. Which inequality shows the number of cakes, x, Shiloh can sell to meet her goal? 20 ≤ x ≤ 200 40 ≤ x ≤ 100 100 ≤ x ≤ 200 20 ≤ x ≤ 100

OpenStudy (anonymous):

sorry taking me to long thinking

OpenStudy (anonymous):

Haha Thanks for trying! (:

OpenStudy (anonymous):

@Wolverine32

OpenStudy (ranga):

Goal is to raise at least $200. At $5 a cake she has to sell a minimum of 200/5 = 40 cakes to meet her goal. The maximum number of cakes she can sell is 100. Therefore, \(40 \le x \le 100\).